The line segment connecting (xsub1,6) and (9.ysub2) is bisected by the point (7,3). Find the values of xsub1, and ysub2.
Answer by Boreal(15235)

The midpoint x value, 7, is the average between the x value of one end and the x value at the other end, (x1+9)/2
Since 7=(x1+9)/2, 14=x1+9, and x1=5
Similarly, (6+y2)/2=3
6+y2=6
y2=0